OpenStudy (anonymous):

Help explain? Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (the_fizicx99):

3x - y = -3 -3x -3x -y = -3x - 3 /-1 /-1 /-1 y = 3x + 3 Y - y1 =m(x-x1) y-(-2)=3(x-4) y + 2 = 3x - 12 -2 -2 y = 3x - 10

OpenStudy (anonymous):

Well first you can write it in slope intercept form. y = mx + b Where m = slope and b = the y intercept. y = 3x + 3 for the equation given. From this, it's given that m, the slope, is 3. Now, we know that for any two lines to be parallel, they must have the same slope. So our new line must have a slope of 3. Now, we know the slope and the given point (4,-2), we can substitute those into the point slope form. y - y1 = m (x - x1) y + 2 = 3 (x - 4) NOW, you would just have to simplify this equation and turn it into slope intercept form once again. y + 2 = 3x - 12 y + 2 - 2 = 3x - 12 -2 y = 3x - 14 This would be the equation that is parallel to 3x - y = - 3. x

OpenStudy (anonymous):

Thank you both. I understand it now. ^^

OpenStudy (the_fizicx99):

Typo, -2 - 12 = -14 >.<